# Weighted Pseudorandom Generators for Read-Once Branching Programs via Weighted Pseudorandom Reductions

### v2

@@ 3.5 Reducing error beyond $1/\operatorname{\mathrm{poly}}(w)$
To further reduce the error from $1/\operatorname{\mathrm{poly}}(w)$ an arbitrary small $\varepsilon$, we use the sampler trick from [Hoz21]. The original technique applies to a PRG and gets a WPRG with a smaller error. Here we also use the technique but we apply it on a WPRG and get a WPRG with a smaller error, for the convenience of our analysis.
We shall first introduce the definition of an averaging sampler:
@@
We now introduce the construction of a WPRG for class $\mathcal{B}_{(n,s,w)}$ with error $\varepsilon$. First, let $(G_{0},w_{0})$ be a WPRG for $\mathcal{B}_{(n,s,w)}$ with error $1/(2w\cdot(n+1)^{2})$ and weight $W$, which is constructed from Lemma 3.12. We then set the parameters as follows:
@@
We assume that $\mathrm{Samp}$ is a $(\alpha,\gamma)$-averaging sampler, and let $K,n_{i,j},w_{i,j}$ be as in Lemma 3.8. Our WPRG is constructed as follows:
$\begin{cases}&G(x,y_{1},\ldots,y_{k},i)=G_{0}(\mathrm{Samp}(x,y_{1}))_{n_{i,1}},G_{0}(\mathrm{Samp}(x,y_{2}))_{n_{i,2}-n_{i,1}},\ldots,G_{0}(\mathrm{Samp}(x,y_{k}))_{n_{i,k}-n_{i,k-1}}\\ &w(x,y_{1},\ldots,y_{k},i)=\sigma_{i}K\cdot\prod_{j=1}^{k}w_{0}(\mathrm{Samp}(x,y_{j}))\\ \end{cases}$
@@
For all $n,s,w$, assume there exists a $W$-bounded $1/(2w\cdot(n+1)^{2})$-WPRG $(G_{0},w_{0})$ for $\mathcal{B}_{(n,s,w)}$ with seed length $d$, and a $(\alpha,\gamma)$-averaging sampler $\mathrm{Samp}:\{0,1\}^{r}\times\{0,1\}^{p}\to\{0,1\}^{d}$ with $\alpha,\gamma$ as defined above. Then there exists a $\varepsilon$-WPRG for $\mathcal{B}_{(n,s,w)}$ with seed length $r+kp$ and weight $W^{\log(n/\varepsilon)/\log(nw)}\cdot\operatorname{\mathrm{poly}}(nw/\varepsilon)$.
@@
$O\left(s+\frac{\log n\log(nw)}{\log\log w-\log\log n}+\log w(\log\log\log w-\log\log\frac{\log w}{\log n/\varepsilon})+\log(1/\varepsilon)\right)$
@@
Let $(G_{0},w_{0})$ be a $1/(2w\cdot(n+1)^{2})$-WPRG for $\mathcal{B}_{(n,s,w)}$ with seed length
$d=O\left(s+\frac{\log n\log(nw)}{\log\log w-\log\log n}+\log w(\log\log\log w-\log\log\frac{\log w}{\log n/\varepsilon})\right)$
and weight $W=\operatorname{\mathrm{poly}}(nw/\varepsilon)$ given by Lemma 3.12.
Let $\mathrm{Samp}$ be a $(1/(2W\cdot w^{2}\cdot(n+1)^{2}),\varepsilon/(2(2n)^{k}\cdot W^{k}))$-averaging sampler with seed length $r=d+O(\log(nw/\varepsilon))$ and $p=O(\log(nw)+\log\log(1/\varepsilon))$ given by Lemma 3.15.
@@
$r+(2k+1)p=O\left(s+\frac{\log n\log(nw)}{\log\log w-\log\log n}+\log w(\log\log\log w-\log\log\frac{\log w}{\log n/\varepsilon})+\log(1/\varepsilon)\right)$
@@ 3.6 Application: a WPRG for regular branching program with a better seed length
Theorem 3.18 (Restate of Theorem 1.7).
For every $w,n\in\mathbb{N}$ with $n\geq\log\log w$, and every $\varepsilon\geq 2^{-n}$, there exists an explicit $\varepsilon$-WPRG for $\mathcal{R}_{n,1,w}$ with seed length
In Section 6.3 of [CHL+23], the WPRG $(G,w):[K]\times(\{0,1\}^{s})^{r}\to\{0,1\}^{n}\times\mathbb{R}$ for $\mathcal{R}_{n,1,w}$ before using the INW generator to replace independent seeds, has the form
@@
$w(i,y_{1},y_{2},\ldots y_{r})$ $=K\cdot\sigma_{i},$
where $i\in[K]$ and $\forall j\in[r],j\in\{0,1\}^{s}$. The parameters satisfy $K=n^{O(m)},r=O(m\log n),\sigma_{i}\in\{-1,+1\}$, where $m=\Theta\left(\frac{\log(w/\varepsilon)}{\sqrt{\log(1/\varepsilon)}+\log w}\right)=O\left(\sqrt{\log(1/\varepsilon)}\right)$. $G_{i,j}$ is the PRG by Braverman, Rao, Raz, and Yehudayoff [BRRY14] for $\mathcal{R}_{n,1,w}$ within target error $\tau=\min\left\{2^{-\sqrt{\log(1/\varepsilon)}},1/\operatorname{\mathrm{poly}}(w,\log n)\right\}$. Therefore, the length of each $y$ is
$s=O\left(\log n\left(\sqrt{\log(1/\varepsilon)}+\log w+\log\log n\right)\right).$
We view the WPRG as a weighted pseudorandom reduction from regular ROBPs to standard ROBPs by fixing $i$ and let $y_{1},\dots,y_{r}$ be free. This is summarized as the following theorem.
Theorem 3.19 (Section 6.3 of [CHL+23]).
For all $n,w\in\mathbb{N}$ and $\varepsilon>0$, there exists a $(\log K,K,\varepsilon)$-weighted pseudorandom reduction from $\mathcal{R}_{n,1,w}$ to $\mathcal{B}_{r,s,w}$, where $K=n^{O(m)},r=O(m\log n)$ and
$s=O\left(\log n\left(\sqrt{\log(1/\varepsilon)}+\log w+\log\log n\right)\right).$
@@
To construct a WPRG that fools $\mathcal{R}_{n,1,w}$ within error $\varepsilon$, we can use a $(\log K,K,\varepsilon/2)$-weighted pseudorandom reduction given by Theorem 3.19 and fool every reduced short ROBP in $\mathcal{B}_{r,s,w}$ within error $\varepsilon/(2K)$. Denote $B$ as the original regular ROBP. Denote $\sum_{i\in[K]}\frac{w_{i}}{K}\cdot\mathbb{E}[B_{i}]$ as the summation of expectations of reduced short ROBPs given by the weighted pseudorandom reduction. Then we have
$\left|\mathbb{E}[B]-\sum_{i\in[K]}\frac{w_{i}}{K}\cdot\mathbb{E}[B_{i}]\right|\leq\varepsilon/2,$
nd $|w_{i}|\leq K$. Denote $(G,w^{\prime}):\{0,1\}^{d}\to(\{0,1\}^{s})^{r}\times\mathbb{R}$ as the $\varepsilon/(2K)$-WPRG. Then we have
$\forall i\in K,\left|\mathbb{E}[B_{i}]-\sum_{x\in\{0,1\}^{d}}\frac{w^{\prime}(x)}{2^{d}}B_{i}(G(x))\right|\leq\varepsilon/(2K),$
Using triangular inequalities, we have
$\left|\mathbb{E}[B]-\sum_{i\in[K]}\frac{w_{i}}{K}\cdot\sum_{x\in\{0,1\}^{d}}\frac{w^{\prime}(x)}{2^{d}}B_{i}(G(x))\right|\leq\varepsilon.$
This gives a WPRG for $B$ with seed length $\log K+d$.
As $r,\log K=O(m\log n)=O\left(\log n\sqrt{\log(1/\varepsilon)}\right)$, using INW generator for fooling $B_{i}$’s would require
$d=s+O\left(\log r\log(Kw/\varepsilon)\right)=s+O\left(\left(\log\log n+\log\log(1/\varepsilon)\right)\cdot\left(\log n\sqrt{\log(1/\varepsilon)}+\log w+\log(1/\varepsilon\right)\right)$
bits of randomness. Instead, our WPRG from Theorem 3.2 for this step requires
$d=O\left(s+\frac{\log r\log w}{\log\log w-\log\log r}+\log w\log\log\log w+\log n\sqrt{\log(1/\varepsilon)}+\log(1/\varepsilon)\right)$
bits.
Note that $\log r=O(\log\log n+\log\log(1/\varepsilon))\leq\log n$ since $\varepsilon\geq 2^{-n}$, and $\log\log\log w\leq\log n$ since $n\geq\log\log w$. Therefore, the overall seed length of our WPRG is
$\log K+O\left(s+\log n\log w+\log(1/\varepsilon)\right)$
$=$ $O\left(\log n\left(\sqrt{\log(1/\varepsilon)}+\log w+\log\log n\right)+\log(1/\varepsilon)\right).$
@@ 4 WPRG for Permutation Read-once Branching Programs
We denote the class of permutation ROBPs with unbounded width and only one accept node by $\mathcal{P}_{(n,s)}$, where $n$ is the length $s$ is the log-size of the alphabet.
Next we show our WPRG for $\mathcal{P}_{(n,s)}$.
@@
Let $(G,w):\{0,1\}^{s}\rightarrow\{0,1\}^{n}\times\mathbb{R}$ be a WPRG. Let $\mathcal{C}$ be a family of matrix valued functions $\mathbf{B}:\{0,1\}^{s}\rightarrow\mathbb{R}^{w\times w}$. We say that $G$ fools $\mathcal{C}$ with $\varepsilon$-sv-error if for all $\mathbf{B}\in\mathcal{C}$, $\sum_{z\in\{0,1\}^{s}}[\frac{1}{2^{s}}\mathbf{B}(G(z))\cdot w(x)]\overset{sv}{\approx}_{\varepsilon}\mathbb{E}_{x\in\{0,1\}^{n}}[\mathbf{B}(x)]$.
We describe a remarkable previous PRG against permutation ROBPs, given by [HPV21b]. From now on we may slightly abuse the notation $\mathcal{P}_{(n,s)}$ to let elements of it also denote matrix-valued functions.
@@
We note that Lemma 4.1 in [CHL+23] only proves the case for $s=2$, but their proof can naturally be generalized to handle arbitrary $s$. See Appendix E.
@@
Then there exists a WPRG $G^{(k)},w^{(k)}$ that can be written in the form of
@@
$w^{(k)}(i)=\sigma(i)\cdot K.$
where $K=n^{O(k)},i\in[K],r=O(k\log n)$, $\sigma(i)\in\{-1,0,1\}$ and $0\leq n_{i,j}\leq n$, such that $G^{(k)}$ fools $\mathcal{P}_{(n,s)}$ with entrywise error $\alpha$.

### v3

@@ 3.5 Reducing error beyond $1/\operatorname{\mathrm{poly}}(w)$
To further reduce the error from $1/\operatorname{\mathrm{poly}}(w)$ an arbitrary small $\varepsilon$, we use the sampler trick from [Hoz21]. The original technique is applied on PRGs to get a WPRG with a smaller error. Here we also use the technique but we apply it on a WPRG and get a WPRG with a smaller error.
We first recall the definition of averaging samplers:
@@
We now give the construction of a WPRG for class $\mathcal{B}_{(n,s,w)}$ with small error $\varepsilon$. First, let $(G_{0},\sigma_{0})$ be a WPRG for $\mathcal{B}_{(n,s,w)}$ with error $1/(2w\cdot(n+1)^{2})$ and weight $W$, which is constructed from Lemma 3.12. We then set the parameters as follows:
@@
We assume that $\mathrm{Samp}$ is a $(\alpha,\gamma)$-averaging sampler, and let $K,n_{i,j},\sigma_{i,j}$ be as in Lemma 3.8. Our WPRG is constructed as follows:
$\begin{cases}&G(x,y_{1},\ldots,y_{k},i)=G_{0}(\mathrm{Samp}(x,y_{1}))_{n_{i,1}},G_{0}(\mathrm{Samp}(x,y_{2}))_{n_{i,2}-n_{i,1}},\ldots,G_{0}(\mathrm{Samp}(x,y_{k}))_{n_{i,k}-n_{i,k-1}}\\ &\sigma(x,y_{1},\ldots,y_{k},i)=\sigma_{i}K\cdot\prod_{j=1}^{k}\sigma_{0}(\mathrm{Samp}(x,y_{j}))\\ \end{cases}$
@@
For all $n,s,w$, assume there exists a $W$-bounded $1/(2w\cdot(n+1)^{2})$-WPRG $(G_{0},\sigma_{0})$ for $\mathcal{B}_{(n,s,w)}$ with seed length $d$, and a $(\alpha,\gamma)$-averaging sampler $\mathrm{Samp}:\{0,1\}^{r}\times\{0,1\}^{p}\to\{0,1\}^{d}$ with $\alpha,\gamma$ as defined above. Then there exists a $\varepsilon$-WPRG for $\mathcal{B}_{(n,s,w)}$ with seed length $r+kp$ and weight $\sigma^{\log(n/\varepsilon)/\log(nw)}\cdot\operatorname{\mathrm{poly}}(nw/\varepsilon)$.
@@
$O\left(s+\frac{\log n\log(nw)}{\log\log w-\log\log n}+\log w(\log\log\min\{n,\log w\}-\log\log\frac{\log w}{\log(n/\varepsilon)})+\log(1/\varepsilon)\right)$
@@
Let $(G_{0},\sigma_{0})$ be a $1/(2w\cdot(n+1)^{2})$-WPRG for $\mathcal{B}_{(n,s,w)}$ with seed length
$d=O\left(s+\frac{\log n\log(nw)}{\log\log w-\log\log n}+\log w(\log\log\min\{n,\log w\}-\log\log\frac{\log w}{\log(n/\varepsilon)})\right)$
and weight $K=\operatorname{\mathrm{poly}}(nw/\varepsilon)$ given by Lemma 3.12.
Let $\mathrm{Samp}$ be a $(1/(2W\cdot w^{2}\cdot(n+1)^{2}),\varepsilon/(2(2n)^{k}\cdot\sigma^{k}))$-averaging sampler with seed length $r=d+O(\log(nw/\varepsilon))$ and $p=O(\log(nw)+\log\log(1/\varepsilon))$ given by Lemma 3.15.
@@
$r+(2k+1)p$
$=$ $O\left(s+\frac{\log n\log(nw)}{\log\log w-\log\log n}+\log w(\log\log\min\{n,\log w\}-\log\log\frac{\log w}{\log(n/\varepsilon)})+\log(1/\varepsilon)\right)$
@@ 3.6 Application: a WPRG for regular branching program with a better seed length
First we recall the ’Weight’ of an ROBP defined in [BRRY14], which is different from the weight of a WPRG.
Definition 3.18 (Weight of ROBPs).
Let $f\in\mathcal{B}_{n,s,w}$ be an ROBP. For every $v\in[n]\times[w]$, let $f^{v\to}$ denote the subsection of $f$ starting at $v$ and has the same accept nodes as $f$. Let $E$ be the edges of $f$. We define the weight of $f$ as:
$\mathsf{W}(f)=\sum_{(u,v)\in E}\left|\mathbb{E}[f^{v\to}]-\mathbb{E}[f^{u\to}]\right|.$
@@
Theorem 3.19.
For every $w,n,s\in\mathbb{N}$ , and every $\varepsilon>0$, there exists an explicit $\varepsilon$-WPRG for $\mathcal{B}_{n,s,w}$ with seed length
$O\left(s+\log n\left(\sqrt{\log(1/\varepsilon)}+\log w+\log(\mathsf{W}(f))+\log\log n\right)+\log(1/\varepsilon)\right).$
Furthermore, the weight of this WPRG is bound by $\operatorname{\mathrm{poly}}(n^{{1+\sqrt{\log(1/\varepsilon)}}},w)$.
One main result of [BRRY14] states that the weights of regular ROBPs over a binary alphabet are bounded by $O(w^{2})$.
Lemma 3.20 (Lemma 6 of [BRRY14]).
For every $n,w\in\mathbb{N}$ and $f\in\mathcal{R}_{n,1,w}$, $\mathsf{W}(f)\leq O(w^{2})$.
Combining Lemma 3.20 with Theorem 3.19 immediately gives Theorem 1.7.
Before we prove Theorem 3.19, we recall the following from [CHL+23, CL24]. In Section 6.3 of [CHL+23], the WPRG $(G,w):[K]\times(\{0,1\}^{s})^{r}\to\{0,1\}^{n}\times\mathbb{R}$ for $\mathcal{R}_{n,1,w}$ , has the form
@@
$\sigma(i,y_{1},y_{2},\ldots y_{r})$ $=K\cdot\gamma_{i},$
where $i\in[K]$ and $\forall j\in[r],j\in\{0,1\}^{s}$. The parameters satisfy $K=n^{O(m)},r=O(m\log n),\gamma_{i}\in\{-1,+1\}$, where $m=\Theta\left(\frac{\log(w/\varepsilon)}{\sqrt{\log(1/\varepsilon)}+\log w}\right)=O\left(\sqrt{\log(1/\varepsilon)}\right)$. $G_{i,j}$ is the PRG by [BRRY14] for $\mathcal{R}_{n,1,w}$ within target error $\tau=\min\left\{2^{-\sqrt{\log(1/\varepsilon)}},1/\operatorname{\mathrm{poly}}(w,\mathsf{W}(f),\log n)\right\}$. Therefore, the length of each $y$ is
$s=O\left(\log n\left(\sqrt{\log(1/\varepsilon)}+\log w+\log(\mathsf{W}(f))+\log\log n\right)\right).$
We view the WPRG as a weighted pseudorandom reduction from regular ROBPs to standard ROBPs by fixing $i$ and let $y_{1},\dots,y_{r}$ be free. Moreover, this WPRG can naturally extend to large alphabet cases although not explicitly stated in [CHL+23]. [CL24] gives a similar result. We restate both of their results as the following theorem, and we include a short proof in Appendix E for completeness.
Theorem 3.21 (Section 6.3 of [CHL+23], also Section 3 of [CL24]).
For all $n,w\in\mathbb{N}$ and $\varepsilon>0$, there exists a $(\log K,K,\varepsilon)$-weighted pseudorandom reduction from $\mathcal{B}_{n,s,w}$ to $\mathcal{B}_{n_{1},n_{1},w}$, where $K=n^{O(\sqrt{\log(1/\varepsilon})},n_{1}=O(\log n\sqrt{\log(1/\varepsilon})$ and
$s_{1}=O\left(s+\log n\left(\sqrt{\log(1/\varepsilon)}+\log w+\log(\mathsf{W}(f))+\log\log n\right)\right).$
@@
In the case of $w>n$, Theorem 3.2 already gives a suitable WPRG. So we assume $w\leq n$. Denote $f\in\mathcal{B}_{n,s,w}$ as the original regular ROBP. Let $(\mathsf{R},\sigma)=\{(\mathsf{R}_{i},\sigma_{i})\}_{i\in[K]}$ be the $(\log K,K,\varepsilon/2)$reduction from $\mathcal{R}_{n,s,w}$ to $\mathcal{B}_{n_{1},s_{1},w}$, given by Theorem 3.21, where $K=n^{O(\sqrt{\log(1/\varepsilon})},n_{1}=O(\log n\sqrt{\log(1/\varepsilon)})$, and
$s_{1}=O\left(s+\log n\left(\sqrt{\log(1/\varepsilon)}+\log w+\log(\mathsf{W}(f))+\log\log n\right)\right).$
Then we have
$\left|\mathbb{E}[f]-\sum_{i\in[K]}\frac{\sigma_{i}}{K}\cdot\mathbb{E}_{x\in(\{0,1\}^{s_{1}})^{n_{1}}}\left[f(\mathsf{R}_{i}(x)\right]\right|\leq\varepsilon/2,$
in which $|\sigma_{i}|\leq K$. Let $(G,\sigma^{\prime}):\{0,1\}^{d}\to(\{0,1\}^{s_{1}})^{n_{1}}\times\mathbb{R}$ be the $\varepsilon/(2K)$-WPRG given by Theorem 3.2. We have
$\left|\mathbb{E}[f]-\sum_{i\in[K]}\frac{\sigma_{i}}{K}\cdot\sum_{y\in\{0,1\}^{d}}\frac{\sigma^{\prime}(y)}{2^{d}}\left[f(\mathsf{R}_{i}(G(y))\right]\right|$
$\leq$ $\left|\mathbb{E}[f]-\sum_{i\in[K]}\frac{\sigma_{i}}{K}\cdot\mathbb{E}_{x\in(\{0,1\}^{s_{1}})^{n_{1}}}\left[f(\mathsf{R}_{i}(x)\right]\right|+\sum_{i\in[K]}\frac{\sigma_{i}}{K}\left|\mathbb{E}_{x\in(\{0,1\}^{s})^{r}}\left[f(\mathsf{R}_{i}(x)\right]-\sum_{y\in\{0,1\}^{d}}\frac{\sigma^{\prime}(y)}{2^{d}}\left[f(\mathsf{R}_{i}(G(y))\right]\right|$
$\leq$ $\varepsilon/2+K\cdot\varepsilon/(2K)=\varepsilon.$
Therefore $(\mathsf{R}\circ G,\sigma\cdot\sigma^{\prime})$ is a $\varepsilon$-WPRG for $f$ with seed length $d_{total}=\log K+d$.
By Theorem 3.2, the $\varepsilon/(2K)$-WPRG requires
$d=$ $O\left(s_{1}+\frac{\log n_{1}\log(n_{1}w)}{\max\left\{1,\log\log w-\log\log n_{1}\right\}}+\log w\log\log\min\{n_{1},\log w\}+\log(K/\varepsilon)\right)$
We use the assumption that $w\leq n$, organize the terms and find
$d_{total}=$ $O\left(s+\log n\left(\sqrt{\log(1/\varepsilon)}+\log w+\log(\mathsf{W}(f))+\log\log n\right)+\log n_{1}\log n+\log(K/\varepsilon)\right)$
$=$ $O\left(s+\log n\left(\sqrt{\log(1/\varepsilon)}+\log w+\log(\mathsf{W}(f))+\log\log n\right)+\log(1/\varepsilon)\right),$
since $\log n_{1}\log(n_{1}w)=O(\log n_{1}\log n),\log w\log\log\min\{n_{1},\log w\}=O(\log n\log n_{1})$.
Now we compute the weight, which is the upper bound of $|\sigma_{i}|\cdot|\sigma^{\prime}(x)|$ over $i\in[K],x\in\{0,1\}^{d}$. By Theorem 3.2, $|\sigma^{\prime}(x)|\leq\operatorname{\mathrm{poly}}(r,w,\varepsilon/2K)=\operatorname{\mathrm{poly}}(n^{{1+\sqrt{\log(1/\varepsilon)}}},w)$. $|\sigma_{i}|\leq K=n^{O(1+\sqrt{\log(1/\varepsilon)})}$. Therefore, the total weight is bound by $\operatorname{\mathrm{poly}}(n^{{1+\sqrt{\log(1/\varepsilon)}}},w)$.
@@ 4 WPRG for Permutation Read-once Branching Programs
We denote the class of permutation ROBPs with unbounded width and only one accept node by $\mathcal{P}_{(n,s)}$, where $n$ is the length $s$ is the bit-length of the alphabet.
Now we give our WPRG for $\mathcal{P}_{(n,s)}$.
@@
[Restatement of Theorem 1.9] There exists an $\varepsilon$-WPRG for $\mathcal{P}_{(n,s)}$ with seed length $s+O(\log n(\log\log n+\sqrt{\log(1/\varepsilon)})+\log(1/\varepsilon))$ and weight $2^{O(\log n\sqrt{\log(1/\varepsilon)})}$.
@@
Let $(G,\sigma):\{0,1\}^{s}\rightarrow\{0,1\}^{n}\times\mathbb{R}$ be a WPRG. Let $\mathcal{C}$ be a family of matrix valued functions $\mathbf{B}:\{0,1\}^{s}\rightarrow\mathbb{R}^{w\times w}$. We say that $G$ fools $\mathcal{C}$ with $\varepsilon$-sv-error if for all $\mathbf{B}\in\mathcal{C}$, $\sum_{z\in\{0,1\}^{s}}[\frac{1}{2^{s}}\mathbf{B}(G(z))\cdot\sigma(x)]\overset{sv}{\approx}_{\varepsilon}\mathbb{E}_{x\in\{0,1\}^{n}}[\mathbf{B}(x)]$.
We describe the PRG against permutation ROBPs from [HPV21b]. In the following descriptions we regard ROBPs of $\mathcal{P}_{(n,s)}$ as matrix-valued functions.
@@
We note that Lemma 4.1 in [CHL+23] only proves the case for $s=2$, but their proof can naturally be generalized to handle arbitrary $s$. See Appendix D.
@@
Then there exists a WPRG $G^{(k)},\sigma^{(k)}$ that can be written in the form of
@@
$\sigma^{(k)}(i)=\gamma_{i}\cdot K.$
where $K=n^{O(k)},i\in[K],r=O(k\log n)$, $\gamma_{i}\in\{-1,0,1\}$ and $0\leq n_{i,j}\leq n$, such that $G^{(k)}$ fools $\mathcal{P}_{(n,s)}$ with entrywise error $\alpha$.